# On Task Vectors and Gradients

### v4

@@ 1 Introduction
Among the various merging techniques, task arithmetic (Ilharco et al., 2022) stands out for its elegant simplicity and surprising empirical effectiveness. This method constructs a “task vector” by taking the weight difference between a finetuned model and its pretrained base, and then builds a multitask model by summing these vectors. Despite widespread usage, a comprehensive theoretical understanding of why and when task arithmetic works has remained elusive. Prior work (Zhou et al., 2025; Lu et al., 2024; Horoi et al., 2025) has suggested that task vectors derived from shorter intervals of finetunings, although less performant in a task-specific sense, are better suited for merging. However, a rigorous explanation for this phenomenon is still lacking.
@@ 2 Task Vectors as Gradients
This section formalizes the relationship between task vectors and multitask loss gradients. For analytical clarity, we model learning as standard full–batch gradient descent with a fixed step size $\eta$, rather than stochastic or mini–batch variants. This assumption, common in theoretical studies of convergence and generalization (Arora et al., 2018; Nikolakakis et al., ), simplifies the derivations while preserving key insights into the learning dynamics.
Vanilla task arithmetic (TA) (Ilharco et al., 2022) constructs a multitask model by:
@@
Let $T$ denote the set of tasks, with size $|T|$. The pretrained model weights are $\theta_{\text{base}}$. For a task $t\in T$, let $\theta_{t}^{(k)}$ be the parameters obtained after $k$ epochs of finetuning on $t$. The task vector is defined as $\tau_{t}^{(k)}:=\theta_{t}^{(k)}-\theta_{\text{base}}$, i.e., the parameter displacement induced by finetuning (Ilharco et al., 2022). Finetuning on task $t$ minimizes the empirical loss
@@ 2.1 Dominance of The First Epoch
Motivated by the strong performance of one-epoch merging, we next examine how much each epoch actually contributes to the overall optimization. Fig. 3(a) plots the epoch-wise normalized gradient norm $\frac{\|\nabla_{\theta}^{(k)}L\|}{\sum_{k^{\prime}=1}^{K}\|\nabla_{\theta}^{(k^{\prime})}L\|}$.We observe that the first epoch contributes the largest share of total gradient norms across training. Even in datasets where this dominance is less pronounced, such as RESISC45 and DTD, we speculate that the initial epoch still largely determines the training direction. As depicted in Fig. 3(b), the gradients from the first $5$ epochs maintain a high cosine similarity (>0.8>0.8) with the first epoch’s gradient. Thus, the effectiveness of TA arises from the alignment between fully trained and first-epoch task vectors, the latter being closer to true task-loss gradients. Fully trained vectors favor task-specific performance, while first-epoch vectors prioritize gradient approximation. Both yield similar multitask performance, but the single-epoch setting is more efficient.
(a) Normalized gradient norms after $5$ epochs of finetuning.
@@ 2.2 Proof Sketch
Stage $1$: Decomposing Single-Task Error
@@
Stage $2$: Error Aggregation across Tasks
@@
2.3 Bounding Second-Order Error Term $C(\ \theta_{\mathrm{MT}}^{(j)}\}\bigr)$
The error term $C(\{\theta_{\mathrm{MT}}^{(j)}\}_{j=1}^{h}\bigr)$ defined in Equation 3 precisely captures the second-order discrepancy between vanilla task arithmetic and true multitask training. To understand how large this discrepancy can become in practical scenarios, we now provide explicit, uniform bounds on the 2-norm of $C$. We achieve these bounds by introducing mild structural assumptions on the neural network model. Specifically, we consider feed-forward networks with bounded weights, bounded inputs, and activations with controlled first and second derivatives. Under these conditions, the error term $C$ can be explicitly bounded, as detailed in the following theorem. Very similar structural assumptions (bounded weights and inputs, Lipschitz- or controlled‐derivative activations) underpin the theoretical analyses in several seminal works (E.g. (Yarotsky, 2017; Bartlett et al., 2017; Golowich et al., 2018)).
@@ 3.1 Empirical Advantage of Premature Task Vectors
Prior work by Zhou et al. (2025) provides experimental evidence consistent with this perspective. In their iterative task arithmetic framework, they start from a pretrained base model, finetune each task for a short interval to obtain task vectors, merge these vectors via task arithmetic to produce a new base model, and repeat this process over multiple rounds. They demonstrate that, under a fixed total finetuning budget, increasing the merge frequency (and thus decreasing the finetuning interval per round) consistently improves multitask performance. This aligns with our claim that early-stage task vectors are more aligned with multitask gradients and therefore yield better merged models.
@@ 3.2 Parameter Space Trajectory
Iterative task arithmetic (Zhou et al., 2025), a setup where models are repeatedly merged using task vectors obtained after minimal finetuning, leads to a gradual evolution of the base model. This serves as a conceptual tool to visualize whether such updates can guide the model toward lower-loss regions. Fig. 4 shows the 2D PCA projection of merged checkpoints produced by different merging strategies. We observe that methods like TIES and Model Breadcrumbs converge to similar basins, while DARE, owing to its stochastic pruning, diverges significantly. The trajectory composed of small merging steps appears to move toward a distinct and lower-loss region.
@@ 3.3 Task Proficiency is not Mergeability
This outcome aligns with the observation from the previous subsection that highly specialized models diverge further in parameter space (Lu et al., 2024), making naive aggregation less effective (see also Fig. 4). From our gradient approximation standpoint, the longer a model finetunes on its task, the more its task vector becomes a "noisy" surrogate for the true multitask gradient. In contrast, lightly finetuned models remain closer to the pretrained base, resulting in smaller, more mergeable updates. This is because, as we have shown, the first epoch’s task vector is an exact scaled negative gradient, and the contributions from subsequent epochs introduce a curvature-controlled second-order error.
These findings aligns with prior work by Ortiz-Jimenez et al. (2024), which attributes the effectiveness of task arithmetic to weight disentanglement, showing that task vectors are more reliable when models remain close to the pretrained initialization, i.e., in the tangent space, where functional components are more linearly separable. Our work offers an orthogonal perspective based on optimization dynamics rather than geometry. Both analyses converge on a similar insight: models that stay closer to the pretrained base are more mergeable. We attribute this advantage to task vectors being more faithful approximations of the task gradients, in contrast to their explanation based on geometric disentanglement or reduced task interference.
@@ 4 Related work
Mode Connectivity and Model Merging: Mode connectivity studies how weights characterize local minima. Frankle et al. (2020) showed linear paths exist between models with shared initialization, and Entezari et al. (2022) demonstrated convergence to a shared basin post-permutation. Permutation-based merging builds on this, with methods like optimal transport matching (Singh and Jaggi, 2020), Git Re-Basin (Ainsworth et al., 2022), and REPAIR (Jordan et al., 2023). More recent approaches explore embedding-space merging (Navon et al., 2023) and cycle-consistent alignment (Crisostomi et al., 2024). Averaging techniques beyond simple averaging (Wortsman et al., 2022), such as population-based merging (Jolicoeur-Martineau et al., 2023), RegMean (Jin et al., 2022), and Fisher-weighted merging (Matena and Raffel, 2021), optimize merging coefficients in the weighted average. Post-merge performance has been shown to suffer from gradient mismatches between individual models (Daheim et al., 2023). Choshen et al. (2022) introduced a method of merging finetuned models to produce an initialization superior to pretraining, leading to improved downstream performance when subsequently finetuned on a target task.
@@
Task Vectors: Task vector methods (Ilharco et al., 2022) represent task-specific updates as deltas from a shared model, enabling arithmetic operations for transfer, forgetting, and composition. Extensions reduce interference via sparsity (Panda et al., 2024), pruning (Yadav et al., 2023; Davari and Belilovsky, 2023), masking (Yu et al., 2023), and optimization (Yang et al., 2023). Some add test-time adaptation via learned modules (Yang et al., 2024) or advocate disentangled finetuning in tangent space (Ortiz-Jimenez et al., 2024). While these works enhance task vectors post-hoc, several works have explored how the duration of finetuning affects model merging. Specifically, Zhou et al. (2025); Lu et al. (2024); Horoi et al. (2025) highlighted the advantage of merging early task vectors that are finetuned for short intervals, yet exhibiting premature task-specific performance. Our work provides the first rigorous theoretical foundation for this phenomenon.
@@
Multitask Learning: Multitask learning (MTL) is a learning paradigm that aims to improve the generalization performance of multiple related tasks by learning them jointly. Early foundational work (Caruana, 1997) established that MTL can enhance performance by leveraging shared representations and inductive biases across tasks. Subsequent studies (Zamir et al., 2018) have demonstrated that identifying and exploiting the structure among visual tasks can lead to significant reductions in labeled data requirements while maintaining performance. However, MTL presents optimization challenges, particularly due to gradient interference between tasks. Methods like Gradient Surgery (Yu et al., 2020) address this by orthogonally projecting gradients to mitigate conflicts. Similarly, Liu et al. (2021) introduces strategies to balance objectives and ensure convergence. Our theory and experiments show task vectors closely align with task gradients, especially in single-epoch finetuning, suggesting that merging such models can effectively approximate multitask learning.
@@ A.2 Proofs of Theorem 2
$K$ is the number of classes and $P$ the number of parameters. Moreover we denote by $J_{\theta}f_{\theta}(x)\in\mathbb{R}^{K\times P}$ the Jacobian of the network with respect to the parameters, $\nabla^{2}_{\theta}f_{\theta}$ is a $3$-dimensional tensor in $\mathbb{R}^{K\times P\times P}$. By chain rule we can obtain that:

### v5

@@ 1 Introduction
Among the various merging techniques, task arithmetic [13] stands out for its elegant simplicity and surprising empirical effectiveness. This method constructs a “task vector” by taking the weight difference between a finetuned model and its pretrained base, and then builds a multitask model by summing these vectors. Despite widespread usage, a comprehensive theoretical understanding of why and when task arithmetic works has remained elusive. Prior work [33, 18, 12] has suggested that task vectors derived from shorter intervals of finetunings, although less performant in a task-specific sense, are better suited for merging. However, a rigorous explanation for this phenomenon is still lacking.
@@ 2 Task Vectors as Gradients
This section formalizes the relationship between task vectors and multitask loss gradients. For analytical clarity, we model learning as standard full–batch gradient descent with a fixed step size $\eta$, rather than stochastic or mini–batch variants. This assumption, common in theoretical studies of convergence and generalization [2, 21], simplifies the derivations while preserving key insights into the learning dynamics.
Vanilla task arithmetic (TA) [13] constructs a multitask model by:
@@
Let $T$ denote the set of tasks, with size $|T|$. The pretrained model weights are $\theta_{\text{base}}$. For a task $t\in T$, let $\theta_{t}^{(k)}$ be the parameters obtained after $k$ epochs of finetuning on $t$. The task vector is defined as $\tau_{t}^{(k)}:=\theta_{t}^{(k)}-\theta_{\text{base}}$, i.e., the parameter displacement induced by finetuning [13]. Finetuning on task $t$ minimizes the empirical loss
@@ 2.1 Dominance of The First Epoch
Motivated by the strong performance of one-epoch merging, we next examine how much each epoch actually contributes to the overall optimization. Fig. 3(a) plots the epoch-wise normalized gradient norm $\frac{\|\nabla_{\theta}^{(k)}L\|}{\sum_{k^{\prime}=1}^{K}\|\nabla_{\theta}^{(k^{\prime})}L\|}$.We observe that the first epoch contributes the largest share of total gradient norms across training. Even in datasets where this dominance is less pronounced, such as RESISC45 and DTD, we speculate that the initial epoch still largely determines the training direction. As depicted in Fig. 3(b), the gradients from the first 55 epochs maintain a high cosine similarity (>0.8>0.8) with the first epoch’s gradient. Thus, the effectiveness of TA arises from the alignment between fully trained and first-epoch task vectors, the latter being closer to true task-loss gradients. Fully trained vectors favor task-specific performance, while first-epoch vectors prioritize gradient approximation. Both yield similar multitask performance, but the single-epoch setting is more efficient.
(a) Normalized gradient norms after 55 epochs of finetuning.
@@ 2.2 Proof Sketch
Stage 11: Decomposing Single-Task Error
@@
Stage 22: Error Aggregation across Tasks
@@
2.3 Bounding Second-Order Error Term $C(\{\ \theta_{\mathrm{MT}}^{(j)}\}\bigr)$
The error term $C(\{\theta_{\mathrm{MT}}^{(j)}\}_{j=1}^{h}\bigr)$ defined in Equation 3 precisely captures the second-order discrepancy between vanilla task arithmetic and true multitask training. To understand how large this discrepancy can become in practical scenarios, we now provide explicit, uniform bounds on the 2-norm of $C$. We achieve these bounds by introducing mild structural assumptions on the neural network model. Specifically, we consider feed-forward networks with bounded weights, bounded inputs, and activations with controlled first and second derivatives. Under these conditions, the error term $C$ can be explicitly bounded, as detailed in the following theorem. Very similar structural assumptions (bounded weights and inputs, Lipschitz- or controlled‐derivative activations) underpin the theoretical analyses in several seminal works (E.g. [29, 3, 11]).
@@ 3.1 Empirical Advantage of Premature Task Vectors
Prior work by Zhou et al. [33] provides experimental evidence consistent with this perspective. In their iterative task arithmetic framework, they start from a pretrained base model, finetune each task for a short interval to obtain task vectors, merge these vectors via task arithmetic to produce a new base model, and repeat this process over multiple rounds. They demonstrate that, under a fixed total finetuning budget, increasing the merge frequency (and thus decreasing the finetuning interval per round) consistently improves multitask performance. This aligns with our claim that early-stage task vectors are more aligned with multitask gradients and therefore yield better merged models.
@@ 3.2 Parameter Space Trajectory
Iterative task arithmetic [33], a setup where models are repeatedly merged using task vectors obtained after minimal finetuning, leads to a gradual evolution of the base model. This serves as a conceptual tool to visualize whether such updates can guide the model toward lower-loss regions. Fig. 4 shows the 2D PCA projection of merged checkpoints produced by different merging strategies. We observe that methods like TIES and Model Breadcrumbs converge to similar basins, while DARE, owing to its stochastic pruning, diverges significantly. The trajectory composed of small merging steps appears to move toward a distinct and lower-loss region.
@@ 3.3 Task Proficiency is not Mergeability
This outcome aligns with the observation from the previous subsection that highly specialized models diverge further in parameter space [18], making naive aggregation less effective (see also Fig. 4). From our gradient approximation standpoint, the longer a model finetunes on its task, the more its task vector becomes a "noisy" surrogate for the true multitask gradient. In contrast, lightly finetuned models remain closer to the pretrained base, resulting in smaller, more mergeable updates. This is because, as we have shown, the first epoch’s task vector is an exact scaled negative gradient, and the contributions from subsequent epochs introduce a curvature-controlled second-order error.
These findings aligns with prior work by Ortiz-Jimenez et al. [22], which attributes the effectiveness of task arithmetic to weight disentanglement, showing that task vectors are more reliable when models remain close to the pretrained initialization, i.e., in the tangent space, where functional components are more linearly separable. Our work offers an orthogonal perspective based on optimization dynamics rather than geometry. Both analyses converge on a similar insight: models that stay closer to the pretrained base are more mergeable. We attribute this advantage to task vectors being more faithful approximations of the task gradients, in contrast to their explanation based on geometric disentanglement or reduced task interference.
@@ 4 Related work
Mode Connectivity and Model Merging: Mode connectivity studies how weights characterize local minima. Frankle et al. [10] showed linear paths exist between models with shared initialization, and Entezari et al. [9] demonstrated convergence to a shared basin post-permutation. Permutation-based merging builds on this, with methods like optimal transport matching [24], Git Re-Basin [1], and REPAIR [16]. More recent approaches explore embedding-space merging [20] and cycle-consistent alignment [6]. Averaging techniques beyond simple averaging [25], such as population-based merging [15], RegMean [14], and Fisher-weighted merging [19], optimize merging coefficients in the weighted average. Post-merge performance has been shown to suffer from gradient mismatches between individual models [7]. Choshen et al. [5] introduced a method of merging finetuned models to produce an initialization superior to pretraining, leading to improved downstream performance when subsequently finetuned on a target task.
@@
Task Vectors: Task vector methods [13] represent task-specific updates as deltas from a shared model, enabling arithmetic operations for transfer, forgetting, and composition. Extensions reduce interference via sparsity [23], pruning [26, 8], masking [30], and optimization [27]. Some add test-time adaptation via learned modules [28] or advocate disentangled finetuning in tangent space [22]. While these works enhance task vectors post-hoc, several works have explored how the duration of finetuning affects model merging. Specifically, Zhou et al. [33], Lu et al. [18], Horoi et al. [12] highlighted the advantage of merging early task vectors that are finetuned for short intervals, yet exhibiting premature task-specific performance. Our work provides the first rigorous theoretical foundation for this phenomenon.
@@
Multitask Learning: Multitask learning (MTL) is a learning paradigm that aims to improve the generalization performance of multiple related tasks by learning them jointly. Early foundational work [4] established that MTL can enhance performance by leveraging shared representations and inductive biases across tasks. Subsequent studies [32] have demonstrated that identifying and exploiting the structure among visual tasks can lead to significant reductions in labeled data requirements while maintaining performance. However, MTL presents optimization challenges, particularly due to gradient interference between tasks. Methods like Gradient Surgery [31] address this by orthogonally projecting gradients to mitigate conflicts. Similarly, Liu et al. [17] introduces strategies to balance objectives and ensure convergence. Our theory and experiments show task vectors closely align with task gradients, especially in single-epoch finetuning, suggesting that merging such models can effectively approximate multitask learning.
@@ A.2 Proofs of Theorem 2
$K$ is the number of classes and $P$ the number of parameters. Moreover we denote by $J_{\theta}f_{\theta}(x)\in\mathbb{R}^{K\times P}$ the Jacobian of the network with respect to the parameters, $\nabla^{2}_{\theta}f_{\theta}$ is a 33-dimensional tensor in $\mathbb{R}^{K\times P\times P}$. By chain rule we can obtain that: